Find the equation of the line containing the point
P(2, 2)
and parallel to the graph of
2x + y = −1.

Answer by Alan3354(69443)

Similar problem.
---------
Parallel lines have the same slope.
=====================================================================
find the equation of the line that passes through (3,-1) and is perpendicular to the line that passes through (-5,2) and (-1,7).
----------
Step 1, find the slope of the line thru the points.
Step 2, lines perpendicular have a slope that's the negative inverse.
step 3, use
-----
y-y1 = m*x-x1) where m = slope and (x1,y1) is the point (3,-1)
-----------
I'll do a similar problem:
**************
find the equation of the line that passes through (2,-2) and is perpendicular to the line that passes through (-1,3) and (-2,-3).
----------
find the slope of the line m1
m1 = diffy/diffx = -6/-1 = 6
Slope of lines perpendicular m2 = -1/m1 = -1/6
*************
y+2 = (-1/6)*(x-2)
PS It's not "the equation," it's an equation.
===========
email via the TY note to check your work.